### v4

@@ 4.2 Ablation Study
Here, we analyze three key points of the proposed method. The first compares local (LoMaR) and global (MAE) mask methods and explains why we apply LoMaR with its modified ViT and MAE’s deep decoder as the backbone. Then, we discuss different target features (LPF, HOG, SAR-HOG, and ours) on the same backbone to show the need for suppressing noise and target features. Finally, we discuss the role of MIM and PGCA in our architecture.
Table 3: Discussion of the MIM framework for SAR-JEPA. MAE performs global masks, and LoMaR performs local masks. LoMaR-SAR is our modified LoMaR with decoder depth, and we use it as our SAR-JEPA backbone. The classification metric is average accuracy (%). Bold indicates the best result, underline indicates the next best result.
@@
    We replace the linear decoder of LoMaR with an eight-layer decoder in MAE to reduce the effect of noise in the reconstructed pixels by increasing the decoder depth. This improvement inspired us to resolve thoroughly SAR image noise with high-quality target features.
Table 4: Results of different target features for SAR-JEPA. LPF: low pass filter. HOG: histogram of oriented gradient. GR: gradient-by-ratio. The baseline is pixel value, which is corrponds to LoMaR-SAR in Table 3. LPF uses the default filter hyperparameters in (Liu et al., 2023). SAR-HOG and two GR method, whose kernel size can be adjusted, use the same multi-scale setting. $\rm{GR_{lin}}$: gradient-by-ratio with linear kernel. $\rm{GR_{Gau}}$: gradient-by-ratio with Gaussian kernel. The classification metric is average accuracy (%). Bold indicates the best result, underline indicates the next best result.
@@
    We finally chose the linear kernel $\rm{GR_{lin}}$ for SAR-JEPA because it needed a lot of time to fine-tune the Gaussian kernel $\rm{GR_{Gau}}$ parameters in pre-training. And we set the Gaussian standard deviation from $\sigma=0.3*((2r+1-1)*0.5-1)+0.8$. It means the standard deviation is related to the kernel size instead of a fixed parameter in Dellinger et al. (2014).
Figure 4: Multi-scale kernel settings for gradient-by-ratio $\rm{GR_{lin}}$. Here, the scale 1/2/3/4 is with $r$ equal to 5/7/13/17, and the multi-scale concat all four scales in feature channel. Multi-scale is more suitable than single scale due to various targets in remote sensing images.
Table 5: Ablation study of our SAR-JEPA, which is a combination of two architectures (MIM and PGCA). The classification metric is average accuracy (%). Bold indicates the best result, underline indicates the next best result.
@@
    MIM is refers to LoMaR-SAR in Table 3, which is also the MIM module of our SAR-JEPA.
@@
    Due to PGCA (Datcu et al., 2023) using complex SAR image’s features as target signal, we follow its ideas and use our target encoder. Here, PGCA is the contrastive learning between deep features and our multi-scale gradient features $\rm{GR_{lin}}$ to test the effectiveness of the target feature as a guided signal, w.r.t, it removes the MIM task in SAR-JEPA.
@@
LoMaR vs. MAE. Since most targets are small objects in remote sensing images, we prefer masking in local areas to learn the contextual relationships near the target better. Compared to MAE, which masks over the whole image, LoMaR randomly selects multiple local squares for masking. In addition to this, LoMaR uses relative positional encoding, input with mask tokens, and linear decoder. As LoMaR emphasizes, the first two are indispensable to prevent feature collapse. When we remove any of them, pre-training in SAR images with LoMaR and LoMaR-SAR have performance degradation and feature collapse.
@@
LoMaR uses a linear decoder to speed up the computation. However, the linear layer is unsuitable for SAR images because SAR image noise is easier than a deep decoder to affect the encoder features through the loss function. In our early experiments, LoMaR performed better than MAE on small batches of high-quality pre-training data. When we extended the pre-training set to different quality datasets, LoMaR’s performance became unstable. This phenomenon inspired us to make improvements to the decoder and target features. As shown in Table 3, LoMaR-SAR, which replaces the linear decoder of LoMaR with an eight-layer decoder in MAE [^4] However, deeper decoders can be detrimental to encoder performance. For example, the twelve-layer decoder is worse than the eight-layer in the MAE paper., has the best result. Therefore, we use LoMaR-SAR as the MIM module for SAR-JEPA. The target encoder for SAR-JEPA is discussed below. It needs unique designs for SAR images and is at the heart of our architecture.
@@
Target feature selection. In Table 4, we compare the results of pixel value, Low Pass Filter (LPF) (Liu et al., 2023), SAR-HOG (Song et al., 2016), and Gradient-by-Ratio (GR) (Dellinger et al., 2014). All the changes are made only in the target features, while other settings remain unchanged. The Pixel in Table 4 is LoMaR-SAR as baseline. The LPF can mitigate noise interference, but we observe a slight drop of LPF compared to Pixel on the MSTAR’s linear probing result. Therefore, inspired by the HOG feature in MaskFeat, we prefer to obtain high-quality target features with suppressing noise.
@@
Although HOG features have shown effectiveness in FG-MAE for SAR scene-level SSL, we found that the HOG features are unsuitable for target-level SSL in Table 4. This result is because speckle noise can cause the gradient computation in small target regions to have many noise points, greatly affecting the quality of target features. Inspired by the physics-guided contrastive architecture using SAR physics features, we investigate the HOG feature variant in the SAR domain, i.e., SAR-HOG. The core of SAR-HOG is the change in the gradient calculation from differential gradient to ratio gradient. As shown in Table 4, we observe a significant improvement compared to HOG due to this modification. However, the sampling method of HOG may lead to the loss of some small target scattering points, such as the discrete points of an aircraft or a small boat. Therefore, we focus on the gradient information, i.e., target edges. Moreover, basic information, such as the shape and edges of an object, is a generic property (Liu et al., 2022; Cheng et al., 2022; Shi et al., 2020).
@@
We consider two versions of GR in SAR. $\rm{GR_{lin}}$ is computing gradient with a linear kerenal (Song et al., 2016) whose $M$ value equals 1. $\rm{GR_{Gau}}$ is computing gradient with a Gaussian kernel (Dellinger et al., 2014) whose $M$ value obeys a Gaussian distribution. Table 4 shows that the two calculation methods have their advantages for different target categories and have similar performance in most cases. The Gaussian kernel can act as a filter, but the parameters need to be adjusted so as not to blur the target edges overly. Because a pre-trained model needs tens of hours to several days, we prefer to use simple and effective methods with scaling data and parameters. Therefore, we use a linear kernel without spending too much time tuning the parameters of the Gaussian kernel at different scales.
@@
Another important thing is that we propose a multi-scale setting instead of single-scale features for various targets included in the pre-training set. Multi-scale can improve the feature representation for various targets in remote sensing. Figure 4 gives the result of different kernel settings. Scale settings have various effects: smaller scales can better extract small target edges, while larger scales are more obvious for noise suppression. Hence, learning target features at different scales is more advantageous than a single scale.
@@
SAR-JEPA is a combination of MIM and PGCA ideas as shown in Figure 1. Specifically, we designed LoMaR-SAR and multi-scale gradient features as a concrete implementation of the MIM and target encoder module. Table 5 demonstrates the equal importance and mutual reinforcement of the two modules. These focus on the pretext task and signal quality of SSL driven by the data. MIM can provide PGCA with a challenging pretext task of learning features with contextual information in the image. PGCA can improve the quality of features by introducing domain knowledge into the embedding of the target encoder. Combining MIM and PGCA, SAR-JEPA learns high-quality contextual features for low-quality noisy SAR data.
@@ 4.3 Comparisons with other methods
Table 6: Fine-grained classification results of different methods. The classification metric is average accuracy (%). Bold indicates the best result, underline indicates the next best result.
@@
Figure 5: Training loss and test accuracy curves for one fine-tuning of MSTAR 10-shot in Table 6. From the figure, effective self-supervised learning decreases the training loss rapidly and converges to a better-generalized region.
As shown in Table 6, we evaluated the SSL methods in few-shot classification tasks. The comparison methods are ImageNet, MAE (He et al., 2022), Feature Guided MAE (FG-MAE) (Wang et al., 2023d), I-JEPA (Assran et al., 2023), LoMaR (Chen et al., 2022a), and Our SAR-JEPA. Table 6 also gives the difference in the mask area and target features. The ImageNet migrated the supervised pre-training weights on the ImageNet as initialization.
Table 6 studies the result of different pre-trained models on fine-grained SAR target classification. Due to the large difference between SAR and RGB images, ImageNet weights are hard to obtain effect features with small samples under fine-tuning, and using linear probing can have a better result with an unsuitable initialization weight. MAE trained on the SAR dataset performs better than ImageNet with fine-tuning, but the speckle noise affects its representation quality. HOG is not an effective solution to the speckle noise because multiplicative noise leads to the virtual dots in the strong scattering region. It results in a performance degradation in FG-MAE with HOG features, and FG-MAE converges slower than MAE in Figure 5. Our method is a special form of the I-JEPA (Assran et al., 2023) method in SAR. But I-JEPA has the feature collapse, in which the learnable predictor fails to learn a good feature space under speckle noise. In Figure 5, I-JEPA converges to a region with poor test performance. Figure 5 also shows that effective self-supervised learning decreases the training loss rapidly and converges to a better-generalized region
Another interesting phenomenon is the range of contexts, where it can be found that LoMaR with local reconstruction can achieve better fine-tuning results in vehicle and ship compared to MAE. This result illustrates the need to consider the range of contextual information in remote-sensing images. However, due to a linear decoder, LoMaR is more susceptible to the effects of speckle noise in the pixels, leading to unstable performance in Table 6 and slow convergence rate in Figure 5.
Although the proposed method achieves better results on the aircraft dataset, it is not as good as vehicle and ship results. In particular, fine-tuning at SAR-ACD 10-shot is worse than the linear probing, indicating slight overfitting phenomena. It is because the aircraft reflects electromagnetic waves with its streamlined structure, resulting in SAR images that may lack the complete aircraft structure, with only some strong scattering points and shadow edges. Besides, our pre-training did not include many aircraft target samples. Because of this, the generic representation in SAR images still needs to be further explored.
@@ 4.4 Visualization
Figure 6: The averaged attention distance of ViT with different methods. The X-axis is the attention head i.e. layer number. MAE focuses on global information, and other methods have local and global attention ranges. It shows that using pixel-level features as a pretext task with a local image scale can enhance the diversity of Vit modeling with SAR images.
Compared to supervised pre-training and contrastive learning, which have a local attention range in the high layer, ViT with MIM has various attention ranges in different layers (Xie et al., 2023b). However, Figure 6 shows that this characteristic of MIM is related to data property and pixel-level pretext task. We first notice that the deep layers of MAE mainly focus on global information due to large SAR image scenes, which differs from MIM’s modeling properties with ImageNet. Then, LoMaR-SAR with local reconstruction acquires the ability to model local information. This phenomenon illustrates that local information is easily lost in remote sensing images, and local reconstruction can better capture contextual details. In addition, PGCA, which removes the mask in SAR-JEPA and has a pixel-level contrast learning task, can acquire local information. It suggests that the property of modeling local and global information originates from the pixel-level pretext task. Finally, our SAR-JEPA with integrated MIM and PGCA can model local and global features. It can be found that the addition of MIM makes SAR-JEPA pay more attention to high-level global contextual information and more diminished attention to local edge features compared to PGCA. Therefore, the differences in the properties between SAR and RGB images make the same self-supervised learning method show different properties, which is a great challenge for the SSL methods in SAR.
@@ 4.5 Scaling Experiment
Figure 7: Scaling experiment of SAR-JEPA in dataset size, model parameters, and training epoch. We find that SAR-JEPA can benefit from these three aspects, but due to the limitation of dataset size, a large model and long training epochs lead to overfitting and reduce the linear probing performance.
MIM can benefit from scaling data, model parameters, and computational resources, but it will suffer from overfitting when using small data and long training epochs on large models (Xie et al., 2023c). Therefore, we evaluate the scalability of SAR-JEPA with linear probing performance on SAR data. Figure 7 gives the scaling experiment in dataset size, model parameters, and training epoch.
Dataset size. A significant performance increase is observed in small-sample classification with increasing data volume. This phenomenon suggests that by extracting high-quality features as guide signals, self-supervised learning performance can be improved by increasing the number of noisy SAR images. Moreover, this result indicates that SSL does not reach saturation in dataset size and has a huge potential for foundation model to be exploited.
Model parameters. As the model parameters increase, we also observe an increase in representation quality. However, Vit-Large suffered from overfitting problems due to insufficient dataset size and has reduced performance in the fine-grained classification of vehicles and aircraft.
Training epoch. Similarly, too-long training epochs result in the overfitting of Vit-Base, and thereby, our default training setting is 200 epochs. Therefore, the next step is to collect more SAR target data to exploit the potential of SSL and foundation models.

### v5

@@ 4.2 Ablation Study
Here, we analyze three key points of the proposed method. The first compares the local (LoMaR) and global (MAE) mask methods and explains why we apply LoMaR with its modified ViT and the deep decoder of MAE as the backbone. Then, we discuss different target features (LPF, HOG, SAR-HOG, and ours) on the same backbone to show the need to suppress noise and target features. Finally, we discuss the role of MIM and PGCA in our architecture.
Table 3: Discussion of the MIM framework for SAR-JEPA. MAE performs global masks, and LoMaR performs local masks. LoMaR-SAR is our modified LoMaR with decoder depth, and we use it as our SAR-JEPA backbone. The classification metric is the average accuracy (%). Bold indicates the best result, and underline indicates the next best result.
@@
    We replace the linear decoder of LoMaR with an eight-layer decoder in MAE to reduce the effect of noise in the reconstructed pixels by increasing the decoder depth. This improvement inspired us to thoroughly resolve SAR image noise with high-quality target features.
Table 4: Results of various target features for SAR-JEPA. LPF: low pass filter. HOG: histogram of oriented gradient. GR: gradient-by-ratio. The baseline is a pixel value that corresponds to LoMaR-SAR, as illustrated in Table 3. LPF uses the default filter hyperparameters (Liu et al., 2023). SAR-HOG and two GR methods, whose kernel size can be adjusted, use the same multi-scale setting. $\rm{GR_{lin}}$: GR with linear kernel. $\rm{GR_{Gau}}$: GR with Gaussian kernel. The classification metric is the average accuracy (%). Bold indicates the best result, and underline indicates the next best result.
@@
    We finally chose the linear kernel $\rm{GR_{lin}}$ for SAR-JEPA because it requires considerable time to fine-tune the Gaussian kernel $\rm{GR_{Gau}}$ parameters in pretraining. we set the Gaussian standard deviation to $\sigma=0.3*((2r+1-1)*0.5-1)+0.8$. This implies that the standard deviation is related to the kernel size instead of a fixed parameter.
Figure 4: Multi-scale kernel settings for $\rm{GR_{lin}}$. Here, the scale 1/2/3/4 has $r$ equal to 5/9/13/17, and the multi-scale concat all four scales in the feature channel. Multi-scale is more suitable than single scale because of its various targets in remote sensing images.
Table 5: Ablation study of our SAR-JEPA combines two architectures (MIM and PGCA). The classification metric is the average accuracy (%). Bold indicates the best result, and underline indicates the next best result.
@@
    MIM is referred to as LoMaR-SAR in Table 3, which is also the MIM module of our SAR-JEPA.
@@
    Due to PGCA (Datcu et al., 2023) using the features of complex SAR image as target signals, we follow its ideas and use our target encoder. Here, PGCA is the contrastive learning between deep features and our multi-scale gradient features $\rm{GR_{lin}}$ to test the effectiveness of the target feature as a guided signal. It removes the MIM task from SAR-JEPA.
@@
LoMaR vs. MAE. Because most targets are small objects in remote sensing images, we prefer masking in local areas to learn the contextual relationships near the target better. Compared with MAE, which masks the entire image, LoMaR randomly selects multiple local squares for masking. In addition, LoMaR uses relative positional encoding, input with mask tokens, and linear decoder. According to LoMaR, the first two are indispensable to prevent feature collapse. When we remove any of them, pretraining in SAR images with LoMaR and LoMaR-SAR has performance degradation and feature collapse.
@@
LoMaR uses a linear decoder to accelerate the computation. However, the linear layer is unsuitable for SAR images because the SAR image noise easily affects the encoder features through the loss function. In our early experiments, LoMaR performed better than MAE on small batches of high-quality pretraining data. When we extended the pretraining set to various quality datasets, the performance of LoMaR became unstable. This phenomenon inspired us to improve the decoder and target features. As shown in Table 3, LoMaR-SAR, which replaces the linear decoder of LoMaR with an eight-layer decoder in MAE [^4] However, deeper decoders can be detrimental to the encoder performance. For example, the 12-layer decoder performs worse than the 8-layer in the MAE paper., exhibits the best result. Therefore, we use LoMaR-SAR as the MIM module for SAR-JEPA. The target encoder for SAR-JEPA is discussed below. It requires unique designs for SAR images and is important to our architecture.
@@
Target feature selection: In Table 4, we compare the results of the pixel value, low pass filter (LPF) (Liu et al., 2023), SAR-HOG (Song et al., 2016), and GR (Dellinger et al., 2014). All the changes are made only in the target features, whereas the other settings remain unchanged. The Pixel shown in Table 4 is LoMaR-SAR used as baseline. The LPF can mitigate noise interference, but we observe a slight drop of LPF compared to Pixel on the linear probing result of MSTAR. Therefore, inspired by the HOG feature in MaskFeat, we prefer to obtain high-quality target features by suppressing noise.
@@
Although HOG features are effective in FG-MAE for SAR scene-level SSL, we observe that HOG features are unsuitable for target-level SSL, as shown in Table 4. This result is because speckle noise can cause the gradient computation in small target regions to have many noise points, thereby significantly affecting the quality of the target features. Inspired by the physics-guided contrastive architecture using SAR physics features, we investigate the HOG feature variant in the SAR domain, that is, SAR-HOG. The core of SAR-HOG is the change in the gradient calculation from the differential gradient to the ratio gradient. As shown in Table 4, we observe a significant improvement compared to HOG owing to this modification. However, the sampling method of HOG may lead to the loss of some small target scattering points, such as the discrete points of an aircraft or a small boat. Therefore, we focus on the gradient information, that is, the target edges. Moreover, basic information, such as the shape and edges of an object, is a generic property (Liu et al., 2022; Cheng et al., 2022; Shi et al., 2020).
@@
We consider two types of GR in SAR. $\rm{GR_{lin}}$ is computes gradient using a linear kernel (Song et al., 2016) whose $M$ value is equal to 1. $\rm{GR_{Gau}}$ computes gradient using a Gaussian kernel (Dellinger et al., 2014) whose $M$ value obeys a Gaussian distribution. Table 4 shows that the two calculation methods have their advantages for various target categories and have similar performance in most cases. The Gaussian kernel can act as a filter, but the parameters must be adjusted to avoid excessive blurring of the target edges. Because a pretrained model requires tens of hours to several days, we prefer to use simple and effective methods with scaling data and parameters. Therefore, we use a linear kernel without spending too much time tuning the parameters of the Gaussian kernel at various scales.
@@
Another important thing is that we propose a multi-scale setting instead of single-scale features for various targets included in the pretraining set. Multi-scale can improve the feature representation of various targets in remote sensing. Figure 4 illustrates the results of the various kernel settings. Scale settings have various effects: smaller scales can better extract small target edges, while larger scales are more obvious to noise suppression. Hence, learning target features at various scales is more advantageous than a single scale.
@@
SAR-JEPA is a combination of MIM and PGCA ideas, as shown in Figure 1. In particular, we designed LoMaR-SAR and multi-scale gradient features as a concrete implementation of the MIM and target encoder module. Table 5 demonstrates the equal importance and mutual reinforcement of the two modules. They focus on the pretext task and signal quality of SSL driven by the data. MIM can provide PGCA using a challenging pretext task of learning features with contextual information in the image. PGCA can improve the quality of the features by introducing domain knowledge into the embedding of the target encoder. Combining MIM and PGCA, SAR-JEPA learns high-quality contextual features for low-quality noisy SAR data.
@@ 4.3 Comparisons with other methods
Table 6: Fine-grained classification results of various methods. The classification metric is the average accuracy (%). Bold indicates the best result, and underline indicates the next best result.
@@
Figure 5: Training loss and test accuracy curves for one fine-tuning of MSTAR 10-shot (Table 6). From the figure, effective SSL decreases the training loss rapidly and converges to a more generalized region.
As shown in Table 6, we evaluated SSL methods in few-shot classification tasks. The comparison methods employed are ImageNet, MAE (He et al., 2022), Feature Guided MAE (FG-MAE) (Wang et al., 2023d), I-JEPA (Assran et al., 2023), LoMaR (Chen et al., 2022a), and our SAR-JEPA. Table 6 illustrates the difference in the mask area and target features. The ImageNet migrated the supervised pretraining weights on the ImageNet for initialization.
Table 6 presents the results of various pretrained models on the fine-grained classification of the SAR target. Owing to the large difference between SAR and RGB images, ImageNet weights cannot easily obtain effective features with small samples when fine-tuning, and using linear probing can demonstrate a better result with an unsuitable initialization weight. MAE trained on the SAR dataset performs better than ImageNet with fine-tuning, but the speckle noise affects its representation quality. HOG is not an effective solution to speckle noise because multiplicative noise leads to virtual dots in the strong-scattering region. This results in a performance degradation in FG-MAE with HOG features, and FG-MAE converges slower than MAE, as shown in Figure 5. Our method is a special form of the I-JEPA (Assran et al., 2023) method in SAR. However, I-JEPA has the feature collapse, in which the learnable predictor fails to learn a good feature space under speckle noise. As depicted in Figure 5, I-JEPA converges to a region with poor test performance. Figure 5 shows that effective SSL rapidly decreases training loss and converges to a more generalized region
Another interesting phenomenon is the range of contexts, where it can be found that LoMaR with local reconstruction can achieve better fine-tuning results on vehicles and ships than MAE. This result illustrates the need to consider the range of contextual information in remote sensing images. However, because of a linear decoder, LoMaR is more susceptible to the effects of speckle noise in pixels, leading to unstable performance illustrated in Table 6 and a slow convergence rate in Figure 5.
Although the proposed method achieves better results on the aircraft dataset, the results are not as good as those of vehicles and ships. In particular, fine-tuning at SAR-ACD 10-shot is worse than the linear probing, indicating slight overfitting phenomena. This is because the aircraft reflects electromagnetic waves with its streamlined structure, resulting in SAR images that may lack the complete aircraft structure, with only some strong-scattering points and shadow edges. Moreover, our pretraining did not include many aircraft target samples. Consequently, the generic representation in SAR images still needs to be further explored.
@@ 4.4 Visualization
Figure 6: The averaged attention distance of ViT using various methods. The X-axis is the attention head, that is, the layer number. Attention distance (Dosovitskiy et al., 2020; Xie et al., 2023b) is computed based on example pictures (we use 16 example pictures) by averaging the distance between the query pixel and all other pixels, weighted by the attention weight (Dosovitskiy et al., 2020). A larger attention distance indicates the ability to focus on the correlation of two more distant pixels in an image. MAE focuses on global information, and other methods focus on local and global attention ranges. This Figure shows that using pixel-level features as a pretext task with a local image scale can enhance the diversity of Vit modeling with SAR images.
Compared to supervised pretraining and contrastive learning, which have a local attention range in the high layer, ViT with MIM has various attention ranges in different layers (Xie et al., 2023b). Figure 6 shows that this characteristic of MIM is related to the data property and the pixel-level pretext task. We first observe that the deep layers of MAE mainly focus on global information owing to large SAR image scenes, which differs from the modeling properties of MIM with ImageNet. Then, LoMaR-SAR with local reconstruction can model the local information. This phenomenon illustrates that local information is easily lost in remote sensing images, and local reconstruction can better capture contextual details. In addition, PGCA, which removes the mask in SAR-JEPA and performs a pixel-level contrast learning task, can obtain local information. This suggests that the property of modeling local and global information originates from the pixel-level pretext task. Finally, our SAR-JEPA with integrated MIM and PGCA can model local and global features. It can be observed that the addition of MIM makes SAR-JEPA concentrate more on high-level global contextual information and more diminished attention to local edge features compared to PGCA. Therefore, the differences in the properties between SAR and RGB images make the same SSL method show various properties, which is a great challenge for the SSL methods in SAR.
@@ 4.5 Scaling Experiment
Figure 7: Scaling experiment of SAR-JEPA in dataset size, model parameters, and training epoch. We observe that SAR-JEPA can benefit from these three viewpoints. However, owing to the limitation of dataset size, a large model and long training epochs lead to overfitting and reduced linear probing performance.
MIM can benefit from scaling data, model parameters, and computational resources, but it suffers from overfitting when using small data and long training epochs in large models (Xie et al., 2023c). Therefore, we evaluate the scalability of SAR-JEPA with linear probing performance on SAR data. Figure 7 illustrates the scaling experiment for the dataset size, model parameters, and training epoch.
Dataset size. A significant performance increase is observed in small-sample classification as data volume increases. This phenomenon suggests that by extracting high-quality features as guide signals, SSL performance can be improved by increasing the number of noisy SAR images. Moreover, this result indicates that SSL does not reach saturation in the dataset size and has a significant potential for the foundation model to be exploited.
Model parameters. As the model parameters increase, we observe an increase in representation quality. However, Vit-Large suffered from overfitting problems because of insufficient dataset size and exhibited reduced performance in the fine-grained classification of vehicles and aircraft.
Training epoch. Similarly, too-long training epochs result in overfitting of Vit-Base; consequently, our default training setting is 200 epochs. Therefore, the next step is to collect more SAR target data to exploit the potential of SSL and foundation models.